OpenStudy (anonymous):

x^2-9x+20 factorise quadratics

OpenStudy (nubeer):

familiar with quadratic formula?

OpenStudy (anonymous):

no :(

OpenStudy (cwrw238):

this one can be factorized

hartnn (hartnn):

can u factorize @Raeee_009 ?

OpenStudy (hba):

Quadratic formula \[x=-b \pm \sqrt{b^2-4ac}/2a\]

OpenStudy (nubeer):

\[x = \frac{ -b \pm \sqrt{b ^{2}-4ac } }{ 2a}\] this is the quadratic formula now look at your equation it's x^2-9x+20 so a =1 ,b = -9 , c = 20 just plug in values you will get 2 values of x

OpenStudy (anonymous):

that is so confusing

OpenStudy (cwrw238):

yes - you asked about factoring quadratics factor to the form (x +/- a)(x +/- b)

OpenStudy (cwrw238):

to find a and b you need the product ad = +20 and sum = -9 can u think of 2 numbers which fit this requirement?

OpenStudy (cwrw238):

* product ab = + 20

OpenStudy (anonymous):

hmm.. still don't get it :(

OpenStudy (cwrw238):

20 = -4 * -5 right?

OpenStudy (anonymous):

yeah

OpenStudy (cwrw238):

and -4 + -5 = ?

OpenStudy (hba):